OpenStudy (anonymous):

I have another question for the same topic. that is: u = (2,k,6), v = (l, 5,3) w = (1,2,3) the question is do there exist scalars k, and l such that the vectors are mutually orthogonal .it means i have to find k, l and u,v,w orthogonal to each other?

OpenStudy (anonymous):

You have to calculate ||u-v||, using the fact that:\[\|v\|=\sqrt{\langle v,v\rangle}\]where <*,*> is the inner product. You also need to use the bilinearity of the inner product. \[\|u-v\|=\sqrt{\langle u-v,u-v\rangle}=\sqrt{\langle u,u\rangle-2\langle u,v\rangle+\langle v,v\rangle} \]Since u and v are orthogonal, <u,v>=0. Since they are unit vectors, <u,u>=1 and <v,v>=1. So you get root 2 on the right hand side of the equation.

OpenStudy (anonymous):

hello joe!!

OpenStudy (anonymous):

just says it is an inner product space, doesn't say what the metric is

OpenStudy (anonymous):

oh nvm i misread your answer sorry

OpenStudy (anonymous):

no no i misread it that is all

OpenStudy (anonymous):

It depends on whether or not we can assume the knowledge that when you have an inner product on a space, it induces a norm. I assumed that was the norm in question >.< my bad.

OpenStudy (anonymous):

you are right, ignore me

OpenStudy (anonymous):

of course it is the norm, i just made a reading error

OpenStudy (anonymous):

They are unit vectors, so their magnitude is \(1\)

OpenStudy (anonymous):

\[\|u+v\|^2=\langle u-v,u-v\rangle=\|u\|^2+\|v\|^2-\langle uv\rangle-\langle v,u\rangle\]

OpenStudy (anonymous):

is it that hard, friends?

OpenStudy (anonymous):

what if i make up a matrix as \[\left[\begin{matrix}2 & k & 6 \\ l & 5 & 3 \\1 & 2 & 3\end{matrix}\right]\] and solve the homogenous system to get the condition, if no solution, so the answer is no. what do you think?

OpenStudy (amistre64):

if they are mutually orthogoanl, does that mean they form a bases in R^3?

OpenStudy (amistre64):

if so, then the determinant would need to NOT be zero

OpenStudy (amistre64):

or, you can setup dots 2 k 6 l 5 3 ------ 2+5k+18=0 l 5 3 1 2 3 ----- L+10+9=0 2 k 6 1 2 3 ----- 2+2k+18 = 0

OpenStudy (amistre64):

i spose that first one id 2L ... l looks like a one to me :)

OpenStudy (anonymous):

ok, change the letter from l to m to be clear. but you are solving the matrix?

OpenStudy (anonymous):

so they are end up at L = -19 and K = 10. ?

OpenStudy (amistre64):

im solving the orthogonal property of vectors

OpenStudy (anonymous):

ok, got it

OpenStudy (amistre64):

2 k 6 l 5 3 ------ 2L+5k+18=0 l 5 3 1 2 3 ----- L+10+9=0; L = -19 2 k 6 1 2 3 ----- 2+2k+18 = 0 ; K = -10 2(-19) + 5(-10) + 18 = -70, so not othogonal

OpenStudy (anonymous):

yeap.!! got it, they can othogonal to (1,2,3) but they are not orthogonal to each other. so , the solution is there not exists any k,l satisfy the problem? am I right?

OpenStudy (anonymous):

thanks a lot

OpenStudy (amistre64):

i believe that is correct :)

OpenStudy (anonymous):

@99smartscore person: no need to get medal, right? sorry for that.

OpenStudy (amistre64):

you can do a grahm schmidt process to find an orthogonal basis but i dont see a way that you can make it out of these 3 vectors :)

OpenStudy (anonymous):

that part is on the previous part of Grahm . so, I am not allowed to use it

OpenStudy (amistre64):

lol, medals are pointless. the thankyou is good enough ;)

OpenStudy (anonymous):

thank you!